OpenStudy (anonymous):

PLEASE HELP!!!!!! I GIVE MEDALS!!!!!!! :) Below are the steps to solve an equation: Step 1: |x - 2| + 3 = 7 Step 2: |x - 2| = 7 - 3 Step 3: |x - 2| = 4 Which of the following is a correct next step to solve the equation? x + 2 = -4 -x - 2 = 4 x + 2 = 4 x - 2 = -4

OpenStudy (the_fizicx99):

You want to get rid of the absolute values, so to do so, you make another equation. |x - 2| = 4 becomes x - 2 = -4

OpenStudy (the_fizicx99):

So you have x - 2 = 4 and x - 2 = -4

OpenStudy (anonymous):

so is the answer the last one?

OpenStudy (the_fizicx99):

yes

OpenStudy (anonymous):

Thanks :)

OpenStudy (the_fizicx99):

yw